OpenStudy (ilovecake):

Which expressions are equivalent to the expression 24 ÷ (12 – 4)? Choose exactly two answers that are correct. A: 18-5 times 3 B: 3 + 6 ÷ 3 C:36 ÷ (10 + 2) D:12 – 4 ÷ 2 – 3

OpenStudy (ilovecake):

whoever answers first will get a medal and I will become a fan

OpenStudy (anonymous):

ok so first we do 12-4 which is 8. 24 divided by 8 is 3. So the answer has to equal the three listed below...

OpenStudy (anonymous):

18-5=13 13*3 does not equal 3 so its not A

OpenStudy (anonymous):

a c

OpenStudy (anonymous):

3+6=9. 9divided by 3 equals 3! so the answer must be B right?

OpenStudy (ilovecake):

I think it is b, c, d.......... right?

OpenStudy (ilovecake):

wait I can't pick more than 2 so it is not b c d

OpenStudy (aum):

@bantaray The multiplication takes precedence over subtraction and so A) is 18-5 times 3 = 18 - 15 = 3

OpenStudy (anonymous):

the answer is B and C @Ilovecake

OpenStudy (ilovecake):

Ok thanks @bantaray

OpenStudy (aum):

The answer is A and C. B: 3 + 6 ÷ 3 Division has to be done before addition so 3 + 6 ÷ 3 = 3 + 2 = 5. Remember the order of operations: PEDMAS P: Parenthesis E: Exponentiation D: Division M: Multiplication A: Addition S: Subtraction

OpenStudy (anonymous):

So @aum, there should be three right answers..

OpenStudy (ilovecake):

Isn't it PEMDAS

OpenStudy (ilovecake):

@bantaray no there can only be 2 right answers................

OpenStudy (aum):

Both are same. D and M have the same level of precedence and so either one can be done before the other. A and S also has the same level of precedence and so either one can be done before the other.

OpenStudy (anonymous):

you're right, I see my mistake! It is A and C.. My bad :)

OpenStudy (aum):

So PEMDAS or PEDMAS both will work because M and D are at the same level of precedence.

OpenStudy (anonymous):

i said a,c